​A 3.70-kg block starts from rest at the top of a 30.09 incline and slides a distance of 1.90 m down the incline in 1.20 s.


A 3.70-kg block starts from rest at the top of a 30.09 incline and slides a distance of 1.90 m down the incline in 1.20 s. 

(a) Find the magnitude of the acceleration of the block. 

(b) Find the coefficient of kinetic friction between block and plane. 

(c) Find the friction force acting on the block. 

(d) Find the speed of the block after it has slid 1.90 m.
Answer #1

a) the magnitude of accleration of the block

from d=1/2 at2

1.90=0.5*a*1.202

a=2.638 m/s2

b) the coffiecient of kinetic friction between block and plane

with no friction the accleration down the block would have been: g sin 30 =4.9 m/s2

\mu =2.638/4.9=0.538

c) the friction force acting on the block and

F(friction)= m(delta a)= 3.70(4.9- 2.638) = 8.3694 N

d) the speed of the block after t has slid 1.90 m

Vt2= V12+2 a (xt-xi) = 0+1.90(2.638)1.90 =3.085m/s

   Vt= 3.085m/s